OpenStudy (anonymous):

The line y=2x+b is tangent to the graph of y= sqrt(x) at the point P = (a, sqrt(a) ). Find P and determine b.

OpenStudy (anonymous):

 \[2x+b=\sqrt{x}\] \[(2x+b)^2=x\] \[4x^2+4xb+b^2=x\] \[4x^2+x(4b-1)+b^2=0\] Discriminant. \[b^2-4ac=0\] \[(4b-1)^2-4(4)(b^2)=0\] \[16b^2-8b+1-16b^2=0\] \[8b=1\] \[b=\frac{1}{8}\]